Journalist Report – March 7th

Sol 23 – Keeping the (space)ship afloat!

“I have good news,” he said. “I have looked at the sky. A rocket is coming to take us all home. It will be here in the early morning.”

– Chapter 23 of The Martian Chronicles by Ray Bradbury

Today Jérémy is writing to you, as Marie must focus on her health. Hopefully, she will be back very soon, and in perfect shape, to narrate our chronicles here much better than I will ever do. In the meantime, Alice and Corentin have accompanied her on her way to the Earth, where she can have surgery.

I must say that it was a solemn moment to see them leave, early in the morning, all equipped with their luggage and their determination. As there were only four of us left at the Mars Desert Research Station, I could not help but think about the pioneers of space exploration. Those who ventured into space alone, or on the surface of the Moon with only one fellow crewmember. Alone in front of the Universe, far away from home.

Quentin and Alexandre mentioned how big the station now seemed, as the four of us were gathered on the Upper Deck. Only then did I fully realize that the operations would become slightly more difficult. Indeed, for safety’s sake, we make sure that at least two of us remain in the Hab at all times. Everyone thus had to change their habits, like Adrien who watered the crops earlier than usual so that he could be at the Hab during this morning’s EVA. It was supposed to be a routine activity, as Quentin and I simply had to change the batteries and retrieve the data of our atmospheric instruments. However, even the easiest tasks feel more complex in this context.

This being said, it was also very interesting for everyone to get the opportunity to experience different roles: As both the Commander (myself) and the Executive Officer (Quentin) were in EVA, we gave the responsibility of the station to Alexandre. With the Health & Safety Officer away from the campus, his duties have been handed over to me. And we all gathered our energy to write the requested daily reports.

Replacing some time dedicated to scientific activities with other tasks often feels a bit frustrating, but that’s how it is: we need to keep all the systems running, while ensuring that no useless risk is taken. Onboard the International Space Station, only half of the schedule is dedicated to science, while the rest consists of operational tasks!

When there is already more than enough to keep a crew of seven people busy, one may think that it must be difficult to share this workload between only four crewmembers. However, I think that we have all learnt a lot from each other and ensured that there were enough redundancies so that any subset of four crewmembers would have managed to keep our basic operations up and running. I must say that I feel very proud about this: it is not a success of some individual crewmembers, but a success of the whole team!

During the past few weeks, we often talked about the fact that, when all gathered around the table, we always had the feeling that there were not enough people and that someone was missing. It is a very strange feeling, but we had to count to make sure that the seven of us were actually there. Now I let you imagine how it feels with only four crewmembers left! As it was already dark and I was climbing the stairs towards the Upper Deck, I almost felt the isolation of the old lighthouse keeper. It is quite romantic to see it this way, but it is only one out of many parallels that you can draw between space and the sea, and it is no coincidence that the European module of the International Space Station is called Columbus.

Like sailors at sea, far from the daily stimuli that the world has to offer, we also feel a bit more affected by some events. For instance, in the afternoon we noticed with pain that the leaves of the tomato crops were starting to turn yellow, as the temperature within the GreenHab was reaching about 40°C/100°F. Even worse: the lifeless body of one of our eight fishes was found in the aquaponics system. In such a restricted group, we quickly get attached to the simplest things. I am not sure whether this event would have had such an effect on the Earth. Here, I can ensure you that we will mourn the passing of this poor creature for quite a while! Adrien has tested the water quality to investigate the potential reason behind this unfortunate event. Nothing seems wrong so far, we will keep you posted.

But please let me finish this report on a positive note. As I am writing this, one of our two moons is magnificently rising in the sky, and I just learnt that the surgery of Marie went smoothly. She woke up from anesthesia and will soon take the shuttle back to the station with the two other crewmembers. Only when you lose something important, you truly understand how valuable it is to you. And God knows we missed all of them during this very long day!

EVA Report – March 7th

Crew 275 EVA Report 07-03-2023

EVA # 19

Author: Alexandre VINAS (Crew Astronomer and HabCom)

Purpose of EVA: Change of batteries for all the atmospheric instruments, as well as data retrieval. Then, we will go to North Ridge with the journalist team in order to showcase a typical exploration scenario, based on week #2 photogrammetry experiment

Start time: 10:36

End time: 12:15


10:36 – Airlock closed

10:46 – EVA Crew leaves the MDRS, with a crew of 2 journalists from the French TV

10:54 – EVA Crew arrives at the destination. All the instruments are going well.

11:15 – Batteries of LOAC and MegaAres changed, data retrieved

11:02 – Retrieving of the data, change of the batteries of PurpleAir

11:18 – Retrieving of the data, and change of the batteries of the field mill

11:23 – EVA crew leaves the instruments and goes to North Ridge in order to showcase the photogrammetry experiment to the journalists

11:40 – EVA crew requests an extra time of 20 minutes to mission support for the EVA. Request approved.

12:05 – EVA crew back to the MDRS

12:15 – End of the depressurization, EVA crew back in the hab

Destination: Between the Hab and Marble Ritual near the road.

Coordinates (use UTM WGS 84): N4251000 E518500

Participants: EVA Leader: Quentin Royer (Crew Engineer), Jeremy Rabineau (Commander)

HAB COM: Alexandre Vinas (Crew Astronomer)

Road(s) and routes per MDRS Map: Drive Cow Dung Road 0110 to a place just before Marble Ritual, then to North Ridge (see map on the attached picture)

Mode of travel: Driving and walking

GreenHab Report – March 7th

GreenHab Officer: Adrien Tison

Environmental control: Heater

Average Temperature: 87.5 °F

Average humidity: 20%

Hours of supplemental light: 0 hour

Daily water usage for crops: 12.68 gallons

Daily water usage for research and/or other purposes: 0 gallons

Water in the blue tank: 73.99 gallons

Time(s) of watering for crops: 8:30 & 15:15

Changes to crops: N/A

Narrative: This morning as it was our crew journalist’s departure for their surgery, we got up pretty early. It also was our second media day. So, I had to water plants and crops earlier.

Unfortunately, when I checked on the fish, I found one stuck to the pump, and thus, dead…

We do not know for sure if this is the reason, but water tests that have been conducted are good, it then shouldn’t be due to water. I will for sure continue testing the water to ensure their safety.

Apart from this unfortunate news, today was quite hot, tomato leaves are turning yellow, so I had to water plants more than usual. I also decided to put stronger sunshades for the day (30% on the left side and 50% on the right side).

Harvest: N/A

Support/supplies needed: N/A

Operations Report – March 7th

Crew 275 Operations Report 07Mar2023
SOL: 23

Name of person filing report: Quentin Royer
Non-nominal systems:
• Toilet
• Spacesuit n°9
• Robotic observatory

Notes on non-nominal systems:

Spirit rover used: No
• Hours: 209.9
• Beginning charge: 100%
• Ending charge: N/A
• Currently charging: Yes

Opportunity rover used: No
• Hours: 113.9
• Beginning charge: 100%
• Ending charge: N/A
• Currently charging: Yes

Curiosity rover used: Yes
• Hours: 221.8
• Beginning charge: 100%
• Ending charge: 70%
• Currently charging: Yes

Perseverance rover used: No
• Hours: 255.9
• Beginning charge: 100%
• Ending charge: N/A
• Currently charging: Yes

General notes on rovers: Nothing to report.

Summary of Hab operations: The Crew conducted routine operations. A team of 2 journalists from the French television was here to film our activities here.
• WATER USE: 10.30
• Water: 355.35 gallons
• Static tank pipe heater: On
• Static tank heater: On
• Toilet tank emptied: Yes

Summary of internet: Starlink (MDRSCampus-Guest) is used during COMMS and during the day for emails and to handle the medical emergency. MDRScampus is used for experimental purpose.

Summary of suits and radios: Spacesuit n°4 is repaired.

Summary of GreenHab operations: The Crew botanist conducted routine operations. A team of 2 journalists filmed inside of the GreenHab.
• WATER USE: 12.68 gallons
• Heater: On
• Supplemental light: No
• Harvest: No

Summary of ScienceDome operations: A team of 2 journalists filmed inside of the ScienceDome.
• Dual split: Heat / From 10 pm to 7 am

Summary of RAM operations: A team of 2 journalists filmed inside of the RAM. Spacesuit n°4 was repaired using the tools recommended by Scott Davis.

Summary of any observatory issues: The robotic observatory is non-nominal.

Summary of health and safety issues: None.

Questions, concerns, and requests to Mission Support:
• In the context of an experiment, we would need to simulate emergency protocols. During the next few days, we will simulate a depressurization at a random time between 10:00 and 18:00, and respond to it as a crew, following a protocol established with researchers from CNES.
In this protocol, we would need to perform an emergency EVA when the alarm rings, to simulate a repair outside of the station. Would it be possible to put on EVA suits, exit by the main airlock and perform an EVA in the immediate vicinity of the station, after a simple radio call to Mission Support?

Sol Summary Report – March 7th

Crew 275 Sol Summary Report 07Mar2023

Sol: 23

Summary Title: Keeping the (space)ship afloat!

Author’s name: Jeremy Rabineau, Commander.

Mission Status: Nominal.

Sol Activity Summary:

With three crewmembers leaving to the hospital early in the morning, only four crewmembers remained at the Hab today. Since we knew that things will be more difficult in this context, especially with the presence of journalists onsite, we planned mainly routine activities.

Due to the late arrival of the journalist, we had to slightly delay the EVA dedicated to changing the batteries of the atmospheric instruments and retrieving their data. As planned, we also took some time to go to North Ridge and shoot nice videos with the journalists, including drone footage. The afternoon was then mainly dedicated to showcasing our different experiments and touring the different buildings of the campus with the journalists.

With the temperature in the GreenHab close to 40°C/100°F, we noticed that the leaves of the tomato crops were turning yellow. We are also very sad to announce that one of our fishes has passed out for an unknown reason. The water tests seem to indicate that this event was not linked to the quality of the water. The GreenHab officer will keep an eye on this in the next few days.

Some news also from our injured crewmember. After another X-ray and an appointment with a surgeon in the end of the morning, they have directly proceeded with surgery in the afternoon. It was a success, and they feel well. Everyone feels relieved to know that the three missing crewmembers will be back to the Hab in a few hours!

Look Ahead Plan:

With the three missing crewmembers back to the Hab in a few hours, we will be able to go back to our nominal planning. No EVA is planned for tomorrow, but we would like to perform an exploration EVA on Thursday (Sol 25) and another EVA dedicated to the disassembling of the atmospheric instruments on Friday (Sol 26).

Anomalies in work: None.

Weather: Sunny, low -3°C/26°F, high 11°C/52°F.

Crew Physical Status: One crewmember injured (broken collarbone / clavicula). Together with the Health & Safety Officer and the Crew Scientist, they left for the hospital in the morning to see a surgeon at the Sevier Valley hospital. The surgery was successfully conducted in the afternoon, and they are about to come back to the MDRS.

EVA: EVA # 19 was successfully conducted, after requesting a delay due to the late arrival of the journalist team from France Televisions. The batteries of the atmospheric instruments have been changed and the data retrieved. Then the EVA crew went to North Ridge to shoot some nice videos with the journalists.

Reports to be filed: Sol Summary, EVA Report, Operations Report, Journalist Report, GreenHab Report, HSO Report.

Support Requested: None.

Copyright © The Mars Society. All rights reserved. | Main Site